OpenStudy (anonymous):

Identify the correct slope and y-intercept of the equation 5x - 3y = -6

OpenStudy (anonymous):

do you know point slope formula?

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (anonymous):

the equation needs to be put into y=mx+b form how would you do that for 5x-3y=-6?

OpenStudy (anonymous):

y=5x+3?

OpenStudy (anonymous):

is the 5 negative or positive?

OpenStudy (anonymous):

positive

OpenStudy (anonymous):

-3y=-5x-6 do you subtract or divide the 3?

OpenStudy (anonymous):

when you brung the 5 over to the right side of the equal sign it was subtracted so it becomes a negative

OpenStudy (anonymous):

oh ok. And would you divide the 3?

OpenStudy (anonymous):

yes y=-5/3-2 does 5/3 and 2 stay negative or positive

OpenStudy (anonymous):

negative

OpenStudy (anonymous):

they turn positive when 2 negatives come together they turn positive \[y=\frac{ 5 }{ 3 }+2\]

OpenStudy (anonymous):

\[y=\frac{ 5 }{ 3 }x+2\]

OpenStudy (anonymous):

so the slope is 5/3 and the y intercept is 2?

OpenStudy (anonymous):

yes